OpenStudy (anonymous):

PLEASE HELP WILL MEDAL AND FAN!!!!!!! What's the answer? –4(7x – 3y – z) – (–6)(9x – z + 2)

OpenStudy (anonymous):

Let's simplify step-by-step. −4(7x−3y−z)−−6(9x−z+2) Distribute: =(−4)(7x)+(−4)(−3y)+(−4)(−z)+54x+−6z+12 =−28x+12y+4z+54x+−6z+12 Combine Like Terms: =−28x+12y+4z+54x+−6z+12 =(−28x+54x)+(12y)+(4z+−6z)+(12) =26x+12y+−2z+12 Answer:=26x+12y−2z+12

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

U welcome